OpenStudy (anonymous):

Please help.. Andrea is designing the seating arrangement for a concert in her local park. To give everyone a good view, each row must have 4 more seats than the row before it, and the first row can only have 10 seats. Help Andrea plan the rest of the seating by solving for how many seats are in row 20. Then explain to Andrea how to create an equation to predict the number of seats in any row. Show your work, and use complete sentences.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Since each row increases by 4 seats, the situation can be modeled by arithmetic sequence : first term = 10 common difference = 4

OpenStudy (anonymous):

So n = 20 right?

ganeshie8 (ganeshie8):

general term of arithmetic sequence can be used to predict the number of seats in any row : \(a_n = 10 + (n-1)*4\)

ganeshie8 (ganeshie8):

Yes ! plugin \(n = 20\) in above equation to get the number of seats in \(20\)th row

OpenStudy (anonymous):

so there would be 86 seats?

ganeshie8 (ganeshie8):

\(\large a_{20} = 10 + (20-1)*4 = 10 + 19*4 = 86\)

ganeshie8 (ganeshie8):

Exactly ! you need to show all the work and explain a bit more okay ?

OpenStudy (anonymous):

Thank you so much! c: I'll be okay, thanks again!

ganeshie8 (ganeshie8):

yw